A disk completed 64.0 revolutions and goes 80 rad/s what is angular accerleration?
1
Expert's answer
2019-02-17T09:35:07-0500

If ω_0=0 rad/s.

ϕ=α t^2/2

ω=αt→t=ω/α

ϕ=α (ω/α)^2/2=ω^2/2α

Angular acceleration:

α=ω^2/2ϕ=〖80〗^2/(2(2π)(60))=8.49 rad/s^2 .